OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? 2x-5y = -5 x + 2y = 11 A. (3, 4) B. (5, 3) C. (7, 2) D. (15, 7)

OpenStudy (anonymous):

@calculusxy @jim_thompson5910 i am pretty sure it is either b or d hold on one sec...

OpenStudy (anonymous):

i will medal and fan!!!!

OpenStudy (anonymous):

i am pretty sure it is b!

OpenStudy (anonymous):

Its B

OpenStudy (anonymous):

with proof plz?

OpenStudy (anonymous):

i wanna know how to do this stuff

Directrix (directrix):

@BarfBag Please show work. I did not get B as the correct option.

OpenStudy (anonymous):

2(5)-5(3)=-5 --> 10-15=-5 -----> -5=-5

OpenStudy (anonymous):

oh wait wait wait wait is this proof?

OpenStudy (anonymous):

2x-5y=-5 | x+2y=11 2x-5y=-5 | 2x+4y=22 Subtract the first equation form the second. 9y=27 y=3 Then substitute. x+2(3)=11 x+6=11 x=5 The final answer is the ordered pair (5,3).

OpenStudy (anonymous):

@Directrix what did u get?

OpenStudy (anonymous):

5+2(3)=11--->5+6=11---->11=11

OpenStudy (anonymous):

therefore, its B

OpenStudy (anonymous):

it sounds right to me, what did u get? @Directrix was it d?

Directrix (directrix):

I found the 3 first and mistakenly labeled it as x. My answer is not correct. @Houdini_Dragon Are you using the Mathpapa.com site?

OpenStudy (anonymous):

oh ok that is alright i have one more question this @BarfBag and @Directrix COULD U HELP WITH THAT?

OpenStudy (anonymous):

sorry caps lock

OpenStudy (anonymous):

So the answer is B, as I stated

Directrix (directrix):

>So the answer is B, as I stated Yes and without work. OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

Well I'm sorry if I offended you, but I inserted the work after I stated the answer.